3. The production function of a firm is described by the following equation 𝑸=𝟏𝟎,πŸŽπŸŽπŸŽπ‘³βˆ’πŸ‘π‘³ 𝟐 where L stands for the units of labor.


a) Draw a graph for this equation. Use the quantity produced in the y-axis, and the units of labor in the x-axis.


b) What is the maximum production level?


c) How many units of labor are needed at that point?


1
Expert's answer
2022-03-24T16:11:49-0400

a) The production curve



b) The Maximum production level

Find the marginal product

"\frac{\Delta Q}{\Delta L}" "=10000-6L"


equate the MP to zero to find L

"10000-6L =0"

"L=1666.667=1667"


The maximum production is therefore;

"Q=10,000(1667)\u22123(1667)^2 =8333333"


c) The units of labour at this point is

1667 units as calculated above